# Mutual Information Free Topological Generalization Bounds via Stability

### v1

@@ Appendix A Trajectory Stability of Projected SGD on Nonconvex Problems
To show the trajectory-stability of SGD, we will first show that it is argument-stable in the sense of Definition 2.1, then we will use Lemma 3.2 to establish the trajectory stability.
To prove the argument stability, we almost exactly follow the steps in [HRS16, Theorem 3.12]. Let us consider a neighboring dataset $S^{\prime}=(z^{\prime}_{1},\dots,z^{\prime}_{n})\in\mathcal{Z}^{n}$ that differs from $S$ by one element, i.e., $z_{i}=z^{\prime}_{i}$ except one element. Consider the SGD recursion on $S^{\prime}$:
@@
[…]
@@
The second parameter is upper-bounded in the proof of [HRS16, Theorem 3.12], which reads as follows:
[…]
@@
Combining Corollary A.1 with our empirical evidence from Section 4.1, we demonstrate the practical validity of our assumption. In the subsequent analysis we provide the reasoning behind our main results, which are based on 3.1.
Appendix B Omitted Proofs
In this section, we present in full details the omitted proofs of our main results in Section 3.
Recent advances on topological generalization bounds are based on a data-dependent Rademacher complexity, leveraging PAC-Bayesian theory on random sets [DVDS24]. In our stability-based approach, the Rademacher complexity also plays a central role. We define the standard Rademacher complexity [SSBD14, BM02] as:
Definition B.1 (Rademacher complexity).
Let $\mathcal{W}$ be a set, $S\in\mathcal{Z}^{n}$, and $\boldsymbol{\epsilon}:=(\epsilon_{1},\dots,\epsilon_{n})\sim\mathcal{U}(\{-1,1\})^{\otimes n}$ be i.i.d. Rademacher random variables (i.e., centered Bernoulli variables). We define:
$\mathbf{Rad}_{S}(\mathcal{W}):=\mathbb{E}_{\boldsymbol{\epsilon}}\left[\sup_{w\in\mathcal{W}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}\ell(w,z_{i})\right].$
B.1 A key technical lemma: Rademacher bound for trajectory-stable algorithms
A key technical element of our approach is the following theorem, which presents an expected Rademacher complexity bound for the worst-case generalization error of trajectory-stable algorithms.
Theorem B.1.
Let $S\sim\mu_{z}^{\otimes n}$ be a dataset. Let $J,K\in\mathbb{N}^{\star}$ be such that $n=JK$, and $\tilde{S}_{J}\in\mathcal{Z}^{J}$ be independent of $S$. Assume that algorithm $\mathcal{A}$ satisfies 3.1 with parameter $\beta_{n}>0$, then:
$\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w))\right]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]+2J\beta_{n}.$
The proof is inspired by the data splitting technique used in [DDS23, Theorem 3.8], which we adapt to our purpose of using the trajectory stability assumption to remove the need for information-theoretic terms.
@@ B.1 A key technical lemma: Rademacher bound for trajectory-stable algorithms
In the following, we consider tree independent datasets $S,S^{\prime},S^{\prime\prime}\sim\mu_{z}^{\otimes n}$, with $S=(z_{1},\dots,z_{n})$ (and similar notation for $S^{\prime}$ and $S^{\prime\prime}$). Up to a slight change in the constant, we can without loss of generality consider $J,K\in\mathbb{N}^{\star}$ such that $n=JK$. We write the index et as a disjoint union as follows:
@@
Given two datasets $S_{1},S_{2}\in\mathcal{Z}^{n}$, we also consider a measurable mapping $w(S_{1},S_{2})$ (whose existence is satisfied in our setting [Mol17]) such that, almost surely:
@@
Recall that $\mathcal{W}_{S}\in\mathrm{CL}(\mathbb{R}^{d})$ (where $\mathrm{CL}(\mathbb{R}^{d})$ denotes the closed subsets of $\mathbb{R}^{d}$) is a data-dependent random trajectory, which we assume is a random compact set [Mol17]. Finally we recall the notation:
@@
This result is based on leveraging our stability assumption to enable the use of the classical symmetrization argument [SSBD14]. Compared to previously proposed modified Rademacher complexity bounds, [FGK+19, SvEH+23], the above theorem has the advantage of directly involving the Rademacher complexity of the data-dependent hypothesis set $\mathcal{W}_{S}$, which tightens the bound and makes it significantly more empirically relevant.
@@
Compared to previous studies [FGK+19, DVDS24], a drawback of our bound compared to these studies is that it only holds in expectation. Moreover, while we directly involve the data-dependent trajectory $\mathcal{W}_{S}$, our Rademacher $\mathbf{Rad}_{\tilde{S}}(\mathcal{W}_{S})$ term uses an independent copy of the dataset, hence, losing some data dependence compared to the term $\mathbf{Rad}_{S}(\mathcal{W}_{S})$ obtained by [DVDS24]. That being said, we show that our techniques still recover certain topological complexities, at the cost of requiring an additional Lipschitz continuity assumption. The Rademacher complexity term appearing in Theorem B.1 uses the data-dependent trajectory $\mathcal{W}_{S}$ and an independent dataset $\tilde{S}$ of size $J$. Following classical arguments [SSBD14, ADS+24], we expect this term to behave as $\mathcal{O}(1/\sqrt{J})$, hence, optimizing the value of $J$ yields a bound of order $\mathcal{O}(\beta_{n}^{1/3})$.
B.2 Proof of Theorems 3.5 and 3.6
@@
Before presenting the proof of Theorems 3.5 and 3.6, we present two technical lemma, whose proof can be found in [ADS+24]. The first lemma relates the Rademacher complexity to the $\alpha$-weighted lifetime sums.
@@
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that 3.4 and 3.3 hold. For any $\alpha\in(0,1]$, we have:
$\mathbf{Rad}_{S}(\mathcal{W})\leq\frac{B}{\sqrt{m}}+B\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}))}{m}},$
@@
This result is a particular case of a bound by [ADS+24] as part of the proof of their Theorem 3.4. The assumptions of [ADS+24, Theorem 3.4] are satisfied by 3.4 and 3.3, and by considering the Euclidean distance of $\mathbb{R}^{d}$ as a pseudometric on $\mathcal{W}$. ∎
The second lemma relates the Rademacher complexity to the positive magnitude introduced by [ADS+24].
@@
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that 3.4 and 3.3 hold. Then, we have for any $\lambda>0$ that:
@@
This result was obtained in the proof of [ADS+24, Theorem 3.5] under a quite general Lipschitz regularity condition. For the sake of completeness, we present the proof in our particular case, as it might be more concrete for some readers.
@@
The proof follows the same strategy as the proof of [ADS+24, lemma B.13], with the difference that we use the definition of magnitude fo compact sets.
It was shown by [Mec15] that a finite set $\mathcal{W}$ has magnitude. More precisely, for any $\lambda>0$ there exists a vector $\beta_{s}:\mathcal{W}\to\mathbb{R}$, such that for any $a\in\mathcal{W}$ we have:
@@
Let $\epsilon=(\epsilon_{1},\dots,\epsilon_{m})\sim\mathcal{U}(\{-1,1\})^{\otimes m}$ be Rademacher random variables. By 3.3 we have for all $a\in\mathcal{W}$:
11 11 $\leq\sum_{b\in\mathcal{W}}e^{-\lambda L\left\|a-b\right\|}\beta_{\lambda L}(b)_{+}$
@@
By 3.4, independence, and Hoeffding’s lemma, we obtain:
@@
We can finally prove Theorems 3.5 and 3.6.
See 3.5
@@
By Theorem B.1, we have (recall that $G_{S}$ has been defined in Equation 3):
@@
with the notations of Theorem B.1 for $\tilde{S}_{J}$. Now we can apply Lemma B.2 to obtain that:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\left(\frac{B}{\sqrt{J}}+B\mathbb{E}\left[\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}_{\left\|\cdot\right\|}(\mathcal{W}_{S,U}))}{J}}\right]\right)+2J\beta_{n},$
@@
See 3.6
Proof.
By Theorem B.1, we have (recall that $G_{S}$ has been defined in Equation 3):
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\mathbb{E}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\lambda L\mathcal{W}_{S,U})\right]+2J\beta_{n},$
with the notations of Theorem B.1 for $\tilde{S}_{J}$. Now we can apply Lemma B.2 to obtain that for any $\lambda>0$, we have:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\left(\frac{\lambda B^{2}}{2J}+\frac{1}{\lambda}\log\left(\mathbf{PMag}(\lambda L\mathcal{W}_{S,U})\right)\right)+2J\beta_{n}.$
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq\frac{\lambda B}{\sqrt{J}}+\frac{2B}{\lambda\sqrt{J}}\log\mathbf{PMag}(\frac{\lambda L\sqrt{J}}{B}\mathcal{W}_{S,U})+2J\beta_{n}.$
B.3 Proof of Lemma 3.2
@@
Consider a data-dependent selection of $\mathcal{A}$, denoted $\omega(\mathcal{W},S)$, in the sense of Definition 3.1. We consider a measurable mapping $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that for any closed set $\mathcal{W}\subset\mathbb{R}^{d}$ and any $w\in\mathbb{R}^{d}$, we have:

### v2

@@ Appendix B Omitted Proofs
To show the trajectory-stability of SGD, we will first show that it is argument-stable in the sense of Definition 2.1, then we will use Lemma 3.2 to establish random set stability.
To prove the argument stability, we almost exactly follow the steps in (Hardt et al., 2016, Theorem 3.12). Let us consider a neighboring dataset $S^{\prime}=(z^{\prime}_{1},\dots,z^{\prime}_{n})\in\mathcal{Z}^{n}$ that differs from $S$ by one element, i.e., $z_{i}=z^{\prime}_{i}$ except one element. Consider the SGD recursion on $S^{\prime}$:
@@
$\mathbb{E}_{U}\left[\|w_{K}-w^{\prime}_{K}\|\right]\leq$ $\mathbb{E}_{U}\left[(\|w_{K}\|+\|w^{\prime}_{K}\|)\>\mathds{1}_{\kappa\leq t_{0}}\right]$
@@
The second parameter is upper-bounded in the proof of (Hardt et al., 2016, Theorem 3.12) (adapted to our setting by noting that the projection on the ball is $1$-Lipschitz), which reads as follows:
$\mathbb{E}_{U}\left[\|w_{K}-w^{\prime}_{K}\|\>\mathds{1}_{\kappa>t_{0}}\right]\leq\frac{2L}{G(n-1)}\left(\frac{K}{t_{0}}\right)^{Gc}.$
@@
B.2 A key technical lemma: Rademacher bound for trajectory-stable algorithms
A key technical element of our approach is the following theorem, which presents an expected Rademacher complexity bound for the worst-case generalization error of trajectory-stable algorithms. Before proceeding to the proof, we recall the following notation for the worst-case generalization error.
$G_{S}(\mathcal{W}):=\sup_{w\in\mathcal{W}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right).$
See 3.4
The proof is inspired by the data splitting technique used in (Dupuis et al., 2023, Theorem 3.8), which we adapt to our purpose of using the trajectory stability assumption to remove the need for information-theoretic terms.
@@ B.2 A key technical lemma: Rademacher bound for trajectory-stable algorithms
In the following, we consider tree independent datasets $S,S^{\prime},S^{\prime\prime}\sim\mu_{z}^{\otimes n}$, with $S=(z_{1},\dots,z_{n})$ (and similar notation for $S^{\prime}$ and $S^{\prime\prime}$). Up to a slight change in the constant, we can without loss of generality consider $J,K\in\mathbb{N}^{\star}$ such that $n=JK$. We write the index set as a disjoint union as follows:
@@
Given two datasets $S_{1},S_{2}\in\mathcal{Z}^{n}$, we also consider a measurable mapping $w(\mathcal{W}_{S_{1}},S_{2})$ (whose existence is satisfied in our setting Molchanov (2017)) such that, almost surely:
@@
Recall that $\mathcal{W}_{S}\in\mathrm{CL}(\mathbb{R}^{d})$ (where $\mathrm{CL}(\mathbb{R}^{d})$ denotes the closed subsets of $\mathbb{R}^{d}$) is a data-dependent random trajectory, which we assume is a random compact set Molchanov (2017). We also denote $\rho_{S}$ the conditional distribution of $\mathcal{W}_{S}$ given $S$. Finally we recall the notation:
@@
This result is based on leveraging our stability assumption to enable the use of the classical symmetrization argument Shalev-Shwartz and Ben-David (2014). Compared to previously proposed modified Rademacher complexity bounds, Foster et al. (2019); Sachs et al. (2023), the above theorem has the advantage of directly involving the Rademacher complexity of the data-dependent hypothesis set $\mathcal{W}_{S}$, which tightens the bound and makes it significantly more empirically relevant.
@@
Compared to previous studies Foster et al. (2019); Dupuis et al. (2024), a drawback of our bound compared to these studies is that it only holds in expectation. Moreover, while we directly involve the data-dependent trajectory $\mathcal{W}_{S}$, our Rademacher $\mathbf{Rad}_{\tilde{S}}(\mathcal{W}_{S})$ term uses an independent copy of the dataset, hence, losing some data dependence compared to the term $\mathbf{Rad}_{S}(\mathcal{W}_{S})$ obtained by Dupuis et al. (2024). That being said, we show that our techniques still recover certain topological complexities, at the cost of requiring an additional Lipschitz continuity assumption. The Rademacher complexity term appearing in Lemma˜3.4 uses the data-dependent trajectory $\mathcal{W}_{S}$ and an independent dataset $\tilde{S}$ of size $J$. Following classical arguments Shalev-Shwartz and Ben-David (2014); Andreeva et al. (2024), we expect this term to behave as $\mathcal{O}(1/\sqrt{J})$, hence, optimizing the value of $J$ yields a bound of order $\mathcal{O}(\beta_{n}^{1/3})$.
B.3 Covering bounds
In this section, we present worst-case generalization bounds over data-dependent random sets based on covering numbers, as they have been defined in Definition˜A.2. These covering bounds serve as a basis to derive some of our other main results, namely, our generalization bounds based on fractal dimension and $\alpha$-weighted lifetime sums. We believe that these additional contributions are also of independent interest.
The following lemma is a consequence of classical covering argument for Rademacher complexity (Shalev-Shwartz and Ben-David, 2014).
Lemma B.1 (Covering lemma).
Let $\mathcal{W}\subset\mathbb{R}^{d}$ be a set with finite diameter. We assume that ˜4.2 holds and that the loss $\ell(\cdot,z)$ is $L$-Lipschitz continuous for all $z\in\mathcal{Z}$. Let $m\in\mathds{N}^{\star}$ and $S\in\mathcal{Z}^{m}$, we have
$\mathbf{Rad}_{S}(\mathcal{W})\leq\inf_{\delta>0}\left(L\delta+B\sqrt{\frac{2\log|N_{\delta}(\mathcal{W})|}{m}}\right)$
The argument is very classical, we reproduce it briefly for the sake of completeness.
Proof.
Let $\delta>0$ and $N:=|N_{\delta}(\mathcal{W})|$ and let $(x_{1},\dots,x_{N})$ be a minimal $\delta$-cover of $\mathcal{W}$. By Lipschitz continuity, we have that:
$\mathbf{Rad}_{S}(\mathcal{W})\leq L\delta+\mathbf{Rad}_{S}(\{x_{1},\dots,x_{N}\}).$
By Massart’s lemma (Shalev-Shwartz and Ben-David, 2014, Lemma 26.8), we have that
$\mathbf{Rad}_{S}(\mathcal{W})\leq L\delta+B\sqrt{\frac{2\log N}{m}}=L\delta+B\sqrt{\frac{2\log|N_{\delta}(\mathcal{W})|}{m}}.$
We conclude by taking the infimum over $\delta>0$. ∎
To illustrate the power of our framework, we present in the next theorem a data-dependent worst-case bounds for stable random sets in terms of covering numbers of $\mathcal{W}_{S,U}$.
Theorem B.2.
Suppose that Assumptions 4.2, 3.1, and 4.1 hold, and that $\mathcal{W}_{S,U}$ is almost surely of finite diameter. Without loss of generality, assume that $\beta_{n}^{-2/3}$ is an integer divisor of nn [^4] Not making this assumption would only lead to minor changes in the bound. We have:
$\mathbb{E}\left[\sup_{w\in\mathcal{W}_{S,U}}\left(\mathcal{R}(w)-\widehat{\mathcal{R}}_{S}(w)\right)\right]\leq 2\mathbb{E}\left[\inf_{\delta>0}\left(\delta L_{S,U}+\beta_{n}^{1/3}\left(1+B\sqrt{2\log|N_{\delta}(\mathcal{W}_{S,U})|}\right)\right)\right],$
where $|N_{\delta}(\mathcal{W}_{S,U})|$ is covering number, i.e., the minimum number of balls of radius $\delta$ that are necessary to cover $\mathcal{W}_{S,U}$.
Proof.
By Lemma˜3.4, we have (recall that $G_{S}$ has been defined in Equation˜4):
with the notations of Lemma˜3.4 for $\tilde{S}_{J}$. By ˜4.1, for fixed $(S,U)$ the loss $\ell(\cdot,\mathcal{Z})$ is $L$-Lipschitz continuous for all $z\in\mathcal{Z}$. Therefore, we can apply Lemma˜B.1 to obtain that
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\mathbb{E}\left[\inf_{\delta>0}\left(\delta L_{S,U}+B\sqrt{\frac{2\log|N_{\delta}(\mathcal{W}_{S,U})|}{J}}\right)\right]+2J\beta_{n},$
Covering bounds are widely used for worst-case bounds over data-independent hypothesis sets (i.e., Example˜1.4) (Shalev-Shwartz and Ben-David, 2014). Theorem˜B.2 extends these results to data-dependent hypothesis sets. Theorem˜B.2 also improves over the data-dependent covering bounds of Dupuis et al. (2024, Corollary 33) by avoiding the use of complicated information-theoretic terms and by allowing the Lipschitz constant to be defined only on the data-dependent set $\mathcal{W}_{S,U}$.
B.4 Proof of Theorem˜4.3
We now present the proof of Theorem˜4.3. This theorem is based on our key technical lemma (Lemma˜3.4), the covering lemma (Lemma˜B.1), and a limit argument that is inspired from previous works (Simsekli et al., 2020; Dupuis and Viallard, 2023).
See 4.3
Proof.
Let $(\Omega,\mathcal{F}_{U},\mathbb{P}_{U})$ denote the probability space to which belongs $U$, with $\mathbb{P}_{U}$ denoting the law of $U$. We also denote $\mathcal{F}$ the $\sigma$-algebra associated with the data space $\mathcal{Z}$. As $\mathcal{W}_{S,U}$ is almost-surely of finite diameter, by definition of the upper box-counting dimension, we have $\mu_{z}^{\otimes n}\otimes\mathbb{P}_{U}$-almost surely that:
$\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U}):=\limsup_{\delta\to 0}\frac{\log|N_{\delta}(\mathcal{W}_{S,U})|}{\log(1/\delta)}=\lim_{\delta\to 0}\sup_{0<r<\delta}\frac{\log|N_{r}(\mathcal{W}_{S,U})|}{\log(1/r)}=:\lim_{\delta\to 0}f_{\delta}(\mathcal{W}_{S,U}).$
Let $(\delta_{k})_{k\geq 0}$ be a decreasing positive sequence in $(0,1)$ with $\delta_{k}\to 0$. Let also $\gamma>0$. By Egoroff’s theorem (Bogachev, 2007), there exists a set $\Omega_{\gamma}\in\mathcal{F}^{\otimes n}\otimes\mathcal{F}_{U}$ such that $\mu_{z}^{\otimes n}\otimes\mathbb{P}_{U}(\Omega_{\gamma})\geq 1-\gamma$, and such that the convergence of $\log f_{\delta_{k}}(\mathcal{W}_{S,U})\to_{k}\log\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U})$ is uniform on $\Omega_{\gamma}$. Therefore, there exists $k_{n,\gamma}>0$ such that for all $k\geq k_{n,\gamma}$, we have for $(S,U)\in\Omega_{\gamma}$
$\log f_{\delta_{k}}(\mathcal{W}_{S,U})\leq\log\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U})+\log 2.$
Therefore, for all $0<\delta<\delta_{k}$, we have by taking the exponential that for $(S,U)\in\Omega_{\gamma}$:
$\log|N_{\delta}(\mathcal{W}_{S,U})|\leq 2\log(1/\delta)\overline{\dim}_{\mathrm{B}}(\mathcal{W}_{S,U}).$
Now we use Lemma˜3.4 to get that for all $J$ we have
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]$ $\leq 2\beta_{n}J+2\mathbb{E}_{S,U,\tilde{S}}\left[\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]$
$\leq 2\beta_{n}J+2B\mathbb{E}_{S,U,\tilde{S}}\left[\mathds{1}_{\overline{\Omega_{\gamma}}}(S,U)\right]+2\mathbb{E}_{S,U,\tilde{S}}\left[\mathds{1}_{\Omega_{\gamma}}(S,U)\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right]$
$\leq 2\beta_{n}J+2B\gamma+2\mathbb{E}_{S,U,\tilde{S}}\left[\mathds{1}_{\Omega_{\gamma}}(S,U)\mathbf{Rad}_{\tilde{S}_{J}}(\mathcal{W}_{S,U})\right].$
We make the choice $J:=\beta_{n}^{-2/3}$, which we assumed for simplicity is an integer dividing $n$, so that it is a valid choice. We conclude the proof by Lemma˜B.1 and setting $\gamma:=1/n$ and $\delta_{n}:=k_{n,1/n}$. ∎
Remark B.1.
The careful reader might notice that the proof above requires some measure-theoretic conditions to hold, in particular, the measurability of the mappings $(S,U)\mapsto\mathbf{Rad}_{S}(\mathcal{W}_{S,U})$ and $(S,U)\mapsto|N_{\delta}(\mathcal{W}_{S,U})|$. These aspects have been extensively studied in existing works on covering / fractal bounds on random sets (Dupuis et al., 2023, 2024). In our paper, we implicitly assume these conditions to be satisfied for the sake of simplicity, and refer to these works for more details. This choice is justified as our main application regards finite trajectories (Example˜1.1), where all this conditions are satisfied.
Remark B.2.
The parameter $\delta_{n}$ appearing in the statement of Theorem˜4.3 might seem intriguing. It can be seen from the proof that this parameters quantifies the uniformity in $n$ of the limit defining the upper box-counting dimension (i.e., if this convergence is uniform in $n$, then $\delta$ is independent of $n$). Such a parameter already appears in (Dupuis et al., 2023) in a similar context and it was shown in (Dupuis et al., 2024) that $\delta$ can be made independent of $n$ under a convergence assumption of the random sets distributions. We refer to these works for further details.
B.5 Proof of Theorem˜4.4
@@ B.3 Covering bounds
Before presenting the proof of Theorem˜4.4, we present two technical lemma, whose proof can be found in Andreeva et al. (2024). The first lemma relates the Rademacher complexity to the $\alpha$-weighted lifetime sums.
@@
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that ˜4.2 and that $\ell(\cdot,z)$ is $L$-Lipschitz continuous on $\mathcal{W}$ for all $z\in\mathcal{W}$. For any $\alpha\in(0,1]$, we have:
$\mathbf{Rad}_{S}(\mathcal{W})\leq\frac{B}{\sqrt{m}}+B\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}(\mathcal{W}))}{m}},$
@@
This result is a particular case of a bound by Andreeva et al. (2024) as part of the proof of their Theorem 3.4. The assumptions of (Andreeva et al., 2024, Theorem 3.4) are satisfied by ˜4.2, the Lipschitz continuity assumption, and by considering the Euclidean distance of $\mathbb{R}^{d}$ as a pseudometric on $\mathcal{W}$. ∎
The second lemma relates the Rademacher complexity to the positive magnitude introduced by Andreeva et al. (2024).
@@
Let $m\in\mathbb{N}^{\star}$, $S\in\mathcal{Z}^{m}$, and $\mathcal{W}\subset\mathbb{R}^{d}$ be a finite set. Assume that ˜4.2 holds and that $\ell(\cdot,z)$ is $L$-Lipschitz continuous on $\mathcal{W}$ for all $z\in\mathcal{W}$. Then, we have for any $\lambda>0$ that:
@@
This result was obtained in the proof of (Andreeva et al., 2024, Theorem 3.5) under a quite general Lipschitz regularity condition. For the sake of completeness, we present the proof in our particular case, as it might be more concrete for some readers.
@@
It was shown by Meckes (2015) that a finite set $\mathcal{W}$ has magnitude. More precisely, for any $\lambda>0$ there exists a vector $\beta_{s}:\mathcal{W}\to\mathbb{R}$, such that for any $a\in\mathcal{W}$ we have:
@@
Let $\epsilon=(\epsilon_{1},\dots,\epsilon_{m})\sim\mathcal{U}(\{-1,1\})^{\otimes m}$ be Rademacher random variables (see Definition˜A.1). By Lipschitz continuity, we have for all $a\in\mathcal{W}$ that
$1$ $\leq\sum_{b\in\mathcal{W}}e^{-\lambda L\left\|a-b\right\|}\beta_{\lambda L}(b)_{+}$
@@
By ˜4.2, independence, and Hoeffding’s lemma, we obtain:
@@
We can finally prove Theorem˜4.4.
See 4.4
@@
Persistent homology bound. By Lemma˜3.4, we have (recall that $G_{S}$ has been defined in Equation˜4):
@@
with the notations of Lemma˜3.4 for $\tilde{S}_{J}$. Now we can apply Lemma˜B.3 to obtain that:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\left(\frac{B}{\sqrt{J}}+B\mathbb{E}\left[\sqrt{\frac{2\log(1+K_{n,\alpha}\mathbf{E}^{\alpha}(\mathcal{W}_{S,U}))}{J}}\right]\right)+2J\beta_{n},$
We make the choice $J:=\beta_{n}^{-2/3}$, which we assumed for simplicity is an integer dividing $n$, so that it is a valid choice. This immediately gives the result.
Positive magnitude bound By Lemma˜3.4, we have (recall that $G_{S}$ has been defined in Equation˜4):
with the notations of Lemma˜3.4 for $\tilde{S}_{J}$. Now we can apply Lemma˜B.4 to obtain that:
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq 2\mathbb{E}\left[\inf_{\lambda>0}\left(\frac{\lambda B^{2}}{2J}+\frac{1}{\lambda}\log\left(\mathbf{PMag}(\lambda L_{S,U}\mathcal{W}_{S,U})\right)\right)+2J\beta_{n}\right].$
$\mathbb{E}\left[G_{S}(\mathcal{W}_{S,U})\right]\leq\mathbb{E}\left[\frac{\lambda B}{\sqrt{J}}+\frac{2B}{\lambda\sqrt{J}}\log\mathbf{PMag}(\frac{\lambda L_{S,U}\sqrt{J}}{B}\mathcal{W}_{S,U})+2J\beta_{n}\right].$
@@
To complete this section, we also show Lemma˜3.2
B.6 Proof of Lemma˜3.2
@@
Consider a data-dependent selection of $\mathcal{A}$, denoted $\omega(\mathcal{W},S)$, in the sense of Definition˜3.1. We consider a measurable mapping $\omega^{\prime}:\mathrm{CL}(\mathbb{R}^{d})\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that for any closed set $\mathcal{W}\subset\mathbb{R}^{d}$ and any $w\in\mathbb{R}^{d}$, we have: